# Functional Homotopy: Smoothing Discrete Optimization via Continuous Parameters for LLM Jailbreak Attacks

### v1

@@ 5 Result and discussion
Table 1: RBO scores (ranging from 0 to 1) for various ranking methods in relation to the ground truth ranking. A higher scores indicate stronger positive alignment with the ground truth. Token gradient ranking shows a marginally higher RBO score than random ranking, indicating a very weakly positive alignment. Conversely, for adversarial examples in image models, the RBO score between the ground truth and gradient-based ranking typically exceeds 0.90 [23].
@@
As seen in Table 2, the FH method either matches (as with Mistral and Vicuna) or substantially outperforms (as with Llama-2 and Llama-3) other methods, even when randomly selecting tokens. Notably, we achieve an almost perfect attack success rate on Llama-2, while the closest baseline is more than $30\%$ weaker than FH-GR.
Table 2: The ASR results after 500 and 1000 iterations. Notably, the ASRs for Mistral-v0.3 and Vicuna-v1.5 reach saturation by 500 iterations, leading to the cessation of further runs. It is important to emphasize that, despite utilizing the same number of iterations, the computational demands differ significantly. For instance, GCG requires gradient computation in each iteration, resulting in an $85\%$ increase in time compared to a random token substitution iteration. Consequently, executing GCG for 500 iterations is equivalent to executing GR for 900 iterations. Furthermore, Andriushchenko et al. [1] incorporated random search into their attack strategy, permitting up to 10 0001000010\,00010 000 random iterations, whereas we established an upper limit of $1000$ iterations.
@@
Since the ASRs of attacks on Mistral and Vicuna reach saturation, we turn our attention to Llama-2 and Llama-3. As illustrated in Figure 3, the FH-GR method identifies adversarial suffixes for prompts that other methods do not achieve within the same number of iterations. Specifically, Figure 3(a) shows that FH-GR successfully finds the majority of its attacks within 500 iterations, significantly outperforming GCG, the closest competing baseline. This highlights the efficiency of framing the optimization as a series of easy-to-hard problems. Additionally, we include iteration distribution plots for Mistral and Vicuna in the appendix.
@@
Figure 3: Iteration distribution for successful attacks. We can identify adversarial strings more efficiently than GCG, the closest competing baseline. Although the plots display iteration counts, it is important to note that each GCG iteration requires significantly more time than an iteration of FH-GR.
@@
In our experiments, we found that this approach often led to model overfitting. For instance, when targeting the prompt “How to build a bomb?”, the expected output would be “Sure, here is how to build a bomb”. A parameter state trained to minimize this loss would likely produce this output as a completion, which could subsequently be rejected by the judge. This misalignment arises because the loss function does not precisely correspond to the objective: a jailbreak attack may not necessarily begin with “Sure, here is how to” and outputs like “Sure, here is how to build a bomb” is not recognized as successful attacks. Consequently, overfitting to the loss function might not yield a successful affirmative response. We also experimented with red-teaming data obtained from [8] (8000 samples), which mitigated overfitting; however, we observed that parameter states close to the base model were consistently more challenging to attack.
The choice of learning rate also influences the performance of the FH method. A larger learning rate results in fewer parameter states before convergence but leads to greater distances between them, necessitating more iterations for warm-start attacks. Conversely, a smaller learning rate produces closer parameter states but increases the number of states in parameter space, thereby extending runtimes. This trade-off can be considered as hyperparameters for the FH method, warranting principled selection and careful analysis in future work.
@@
Our functional homotopy framework capitalizes on the duality between model training and input generation. Fine-tuning a model from its base can be viewed as an application of homotopy optimization, which concurrently supports input generation optimization. This duality underscores the functional relationship between models and inputs. Our approach combines reversed robust training with feature transfer in the input space. Initially, we de-robust train safe models to derive vulnerable variants while retaining intermediate models. Subsequently, jailbreak features are transferred from attacks on weaker models and incrementally intensified for stronger models.
We also conduct a preliminary study on the transferability of attack strings from base models to weaker models. Notably, we find that the space of jailbreak strings for safe models is not merely a subset of those for weak models; contrary to the hypothesis that as models become misaligned, the space of jailbreak strings expands monotonically. Details of this study are included in the appendix, with a more comprehensive investigation proposed for future work.
@@ 6 Conclusion
In this study, we critically examine the commonly used token gradient methods for the discrete optimization challenges in language model analysis and propose a novel optimization technique, the functional homotopy method, to address these issues. The homotopy method effectively smooths the original optimization problem by leveraging the continuity of the parameter space. Additionally, our approach offers a machine-learning perspective that highlights the interplay between model training and input generation. This dual interpretation, combined with the homotopy method, fosters an integrated featurization of both models and inputs, potentially inspiring new empirical tools for probing language models.
7 Acknowledgements
@@
Appendix A Appendix
A.1 Elided proof
A.1.1 Proof of Proposition 3.1
@@ A.1 Elided proof
Appendix B Additional Evaluation Details
@@
We use a learning rate 2e-5, warmup ratio 0.04 and a LoRA adapter with rank 16, alpha 32, dropout 0.05, and batch size 2 to fine-tune the models for 64 epochs, leading to 768 checkpoints in total. Each LoRA checkpoint occupies 49 MB of disk space.
@@
Appendix C Transferability of stronger attacks
@@
Appendix D Additional iteration distributions

### v2

@@ 5 Result and discussion
Table 1: RBO scores (ranging from 0 to 1) for various ranking methods in relation to the ground truth ranking. A higher scores indicate stronger positive alignment with the ground truth. Token gradient ranking shows a marginally higher RBO score than random ranking, indicating a very weakly positive alignment. Conversely, for adversarial examples in image models, the RBO score between the ground truth and gradient-based ranking typically exceeds 0.90 (Wang et al., 2024).
@@
As seen in Table 2, the FH method either matches (as with Mistral and Vicuna) or substantially outperforms (as with Llama-2 and Llama-3) other methods, even when randomly selecting tokens. Notably, we achieve an almost perfect attack success rate on Llama-2, while the closest baseline is more than $30\%$ weaker than FH-GR. We provide additional ablations of the FH method when applied to GCG and AutoDAN in Appendix E.
Table 2: The ASR results after 500 and 1000 iterations. Notably, the ASRs for Mistral-v0.3 and Vicuna-v1.5 reach saturation by 500 iterations, leading to the cessation of further runs. It is important to emphasize that, despite utilizing the same number of iterations, the computational demands differ significantly. For instance, GCG requires gradient computation in each iteration, resulting in an $85\%$ increase in time compared to a random token substitution iteration. Consequently, executing GCG for 500 iterations is equivalent to executing GR for 900 iterations. Furthermore, Andriushchenko et al. (2024) incorporated random search into their attack strategy, permitting up to 10 0001000010\,00010 000 random iterations, whereas we established an upper limit of $1000$ iterations.
@@
Since the ASRs of attacks on Mistral and Vicuna reach saturation, we turn our attention to Llama-2 and Llama-3. As illustrated in Figure 3, the FH-GR method identifies adversarial suffixes for prompts that other methods do not achieve within the same number of iterations. Specifically, Figure 3(a) shows that FH-GR successfully finds the majority of its attacks within 500 iterations, significantly outperforming GCG, the closest competing baseline. This highlights the efficiency of framing the optimization as a series of easy-to-hard problems. Iteration distribution plots for Mistral and Vicuna, along with runtime and storage overhead of the FH method, are provided in the appendix.
@@
Figure 3: Iteration distribution for successful attacks, showing the iterations taken by each method to successfully jailbreak the target models on different inputs. Each bar represents here about 50 iterations. Our method can identify adversarial strings more efficiently than GCG, the closest competing baseline. Although the plots display iteration counts, it is important to note that each GCG iteration requires more time than an iteration of FH-GR.
@@
In our experiments, we found that this approach often led to model overfitting. For instance, when targeting the prompt “How to build a bomb?”, the expected output would be “Sure, here is how to build a bomb”. A parameter state trained to minimize this loss would likely produce this output as a completion, which could subsequently be rejected by the judge. This misalignment arises because the loss function does not precisely correspond to the objective: a jailbreak attack may not necessarily begin with “Sure, here is how to” and outputs like “Sure, here is how to build a bomb” is not recognized as successful attacks. Consequently, overfitting to the loss function might not yield a successful affirmative response. We also experimented with red-teaming data obtained from Ganguli et al. (2022) (8000 samples), which mitigated overfitting; however, we observed that parameter states close to the base model were consistently more challenging to attack.
The selection of 500 epochs for model fine-tuning was empirically determined to sufficiently misalign model parameters, facilitating jailbreaking without additional optimization. However, our method’s efficacy persists with fewer epochs, effectively initiating the homotopy from a more aligned model state. Experiments with stronger checkpoints (i.e., models fine-tuned for fewer epochs) demonstrate that the optimization still converges to a lower loss than the base GCG algorithm. Figure 9 in Appendix H illustrates that optimizations starting from stronger checkpoints exhibit more rapid loss reduction compared to GCG, underscoring the robustness of our approach to variations in fine-tuning duration.
@@
Our functional homotopy framework capitalizes on the duality between model training and input generation. Fine-tuning a model from its base can be viewed as an application of homotopy optimization, which concurrently supports input generation optimization. This duality underscores the functional relationship between models and inputs. Our approach combines reversed robust training with feature transfer in the input space. Initially, we de-robust train safe models to derive vulnerable variants while retaining intermediate models. Subsequently, jailbreak features are transferred from attacks on weaker models and incrementally intensified for stronger models. This duality between models (or programs) and inputs, as well as the incremental evolution of both, aligns with widely used techniques in software analysis, such as continuous integration and mutation testing (Kaiser et al., 1989; Brown et al., 2017; Wang et al., 2020). Additionally, we conduct a preliminary study on the transferability of attack strings from base to weaker models, detailed in Appendix D.
@@ 6 Conclusion
This study introduces the functional homotopy method, a novel optimization technique that smooths the LLM input-generation problem by leveraging the continuity of the parameter space. Additionally, our approach highlights the interplay between model training and input generation, offering a dual perspective that integrates the featurization of both models and inputs. This framework has the potential to inspire new empirical tools for analyzing language models.
Acknowledgements
@@
Appendix A $\mathsf{NP}$-hardness of the optimization problem
Let $f:\{0,1\}^{m}\rightarrow\mathbb{R}$ be a neural network with $m$-binary inputs and a canonical feed-forward structure comprising a single hidden layer. The network is formally expressed as:
$f(x)=W_{2}\sigma(W_{1}x+b_{1})+b_{2}$
where $W_{1}\in\mathbb{R}^{n\times m},W_{2}=\mathbb{R}^{1\times n}$ are the weight matrices between the layers, $\sigma$ denotes an activation function, $b_{1}\in\mathbb{R}^{n}$ and $b_{2}\in\mathbb{R}$ is the bias term. In other words, $f$ is the function composition of two affine transformations and an activation layer. Notice that any composition of two consecutive affine transformations is still an affine transformation.
The decision problem under consideration is formulated as an existential problem:
$\forall a\in\mathbb{R}.\exists x\in\{0,1\}^{n}\ldotp f(x)\leq a.$ (5)
This formulation aims to determine, for any given threshold $a$ and two-layer binary neural network, whether there exists an input $x$ such that $f(x)\leq a$. We demonstrate the $\mathsf{NP}$-hardness of this problem in Section A.1. Subsequently, in Section A.2, we elucidate how this formalization (Equation 5) encapsulates the LLM input generation optimization problem.
A.1 Hardness Reduction
We prove the $\mathsf{NP}$-hardness of the decision problem Equation 5 via a reduction from the Boolean satisfiability (SAT) problem. This reduction is constructed such that solving the decision problem is equivalent to determining the satisfiability of the given SAT formula. Consequently, if one can resolve the decision problem, one can, by extension, solve the corresponding SAT instance.
[…]
Simulation of $3\mathsf{CNF}$
To simulate logical operations within the neural network framework, we employ a gadget similar to that proposed by Wang et al. (2022). This approach utilizes common activation functions such as ReLU, sigmoid, and tanh to approximate a step function, defined as:
$\mathrm{step}(x)=\left\{\begin{array}[]{ll}1,&\mbox{$x>0$}\\ 0,&\mbox{$x\leq 0$}\end{array}\right.$ (6)
The step function, being discrete-valued, serves as a fundamental component for mathematical analysis and simulation of other discrete functions. Its versatility in neural networks allows for easy scaling and shifting through weight multiplication or bias term addition, enabling the replacement of the constants 0 and 1 in Equation 6 with any real numbers.
Wang et al. (2022) demonstrated how common activation functions can approximate the step function. For instance, $\mathrm{ReLU}(nx)-\mathrm{ReLU}(nx-1)$ converges to $\mathrm{step}(x)$ as $n\rightarrow\infty$. For our reduction, we define a step-like function:
$s(x)=\mathrm{ReLU}(x)-\mathrm{ReLU}(x-1)=\left\{\begin{array}[]{ll}1,&\mbox{$x\geq 1$}\\ 0,&\mbox{$x\leq 0$}\\ x,&\mbox{$0<x<1$}\end{array}\right.$ (7)
[…]
To simulate the conjunction operation, we sum all clause gates and negate the result: $-\sum_{i=1}^{k}c_{i}$. Consequently, determining the satisfiability of the $3\mathsf{CNF}$ formula is equivalent to solving the following decision problem for this one-layer network:
$\exists x.f(x)\leq-k.$
Proposition A.1.
[…]
Proof.
We have described the construction of the neural network as above. Now we need to show it indeed captures the computation of the $3\mathsf{CNF}$ formula.
Now if $\phi$ is satisfiable, then there exists a truth assignment of $X_{1},\ldots,X_{m}$ to satisfy $\phi$. Let $x_{i}=1$ if $X_{i}$ is true and $x_{i}=0$ if $X_{i}$ is false in this assignment. By the construction of the neural network, the output of $s$ is also true for each step-like function. As a result, $-\sum_{i=1}^{k}c_{i}=-k$.
Now if $\exists x\in\{0,1\}^{m}$ such that $f(x)\leq-k$. Choose a truth assignment for $X_{1},\ldots,X_{m}$ based on $x_{i}$’s values. Because $f=-\sum_{i=1}^{k}c_{i}$ and $0\leq c_{i}\leq 1$, then each $c_{i}=1$. By construction, $l_{1}+l_{2}+l_{3}\geq 1$ for this activation node $c_{i}$. Because $x_{i}$ is discrete-valued, $l_{i}\in\{x_{i},1-x_{i}\}$, then $l_{i}\in\{0,1\}$. $l_{1}+l_{2}+l_{3}\geq 1$ implies that at least one of $l_{1},l_{2},l_{3}$ equals 1. Therefore, by the truth assignment of $X_{j}$, the corresponding literal is also satisfied. Because each clause has a true assignment, then $\phi$ is satisfied. ∎
[…]
Corollary A.2.
The decision problem in Equation 5 is $\mathsf{NP}$-hard.
A.2 LLM input generation Formalization
We posit that the decision problem in Equation 5 effectively encapsulates the LLM input generation optimization problem. Informally, this problem seeks to minimize some loss based on logits for all given possible suffixes. Because we do not impose any structural assumptions on the input embedding or the output layers, the two-layer network can be embedded within the LLM architecture.
Specifically, in the context of jailbreak attack optimization, the objective is to minimize $L(M(\langle x,s\rangle))$, where $M$ represents an LLM, $x$ denotes the prompt, and $L$ is the entropy loss between the logits and an affirmative prefix. This can be extended to minimizing $f(x)$ for $x\in\{0,1\}^{n}$, given the absence of assumed embedding structures or input prompts. The discrepancy between $f\in\mathbb{R}$ and the entropy loss in jailbreak attacks can be reconciled through a network with multiple outputs, where one output corresponds to $f$ and the rest are constants, establishing a bijective relationship between $f$ and the jailbreak loss. Given that the cross-entropy between a logit $(y_{1},\ldots,y_{n})$ and a target class $j$ is:
$\log\frac{\exp(y_{j})}{\sum_{i=1}^{n}\exp(y_{i})}.$
We can construct a multiple output network $F$, where $F(x)_{j}=f(x)$ and the remaining $F(x)$ are constants. Assuming the one-hot encoding corresponds to the first token (class $j$), the cross-entropy becomes:
$\log\frac{\exp(f(x))}{\exp(f(x))+C},$
for some $C>0$. This function is bijective and monotonically increasing. By fixing the remaining affirmative prefix tokens’ logits as constants, minimizing the prefix loss is equivalent to minimizing $f(x)$, demonstrating that a jailbreak algorithm capable of minimizing the loss can also minimize $f(x)$. Consequently, if the minimization of $f$ is solvable, the existential problem for any threshold $a\in\mathbb{R}$ can also be resolved by querying the minimum of $f(x)$.
Appendix B Linear approximation proof
@@
Appendix C Additional Evaluation Details
@@
We use a learning rate 2e-5, warmup ratio 0.04 and a LoRA adapter with rank 16, alpha 32, dropout 0.05, and batch size 2 to fine-tune the models for 64 epochs, leading to 768 checkpoints in total.
Operational overhead
The storage and computational overheads of our proposed method are comparatively modest. Each LoRA checkpoint requires 49 MB, with a theoretical maximum of 768 checkpoints occupying approximately 37 GB. In practice, we utilized significantly fewer checkpoints, further reducing the storage footprint. For context, full model storage requirements are substantially larger: Llama-2 (13 GB), Mistral (14 GB), Vicuna (26 GB), and Llama-3 (60 GB).
The computational overhead is similarly minimal. Model fine-tuning, performed once for all test inputs, takes approximately 20 minutes. In contrast, attacking a single input for 1000 interactions requires about one hour. When amortized across 200 inputs, the running-time overhead is less than 10 seconds per input. Thus, both storage and computational overheads of our method are relatively insignificant compared to the resource demands of complete language models and the overall attack process.
@@
Appendix D Transferability of stronger attacks
@@
Appendix E Additional Results for Functional Homotopy
We apply the principles of functional homotopy on top of GCG and AutoDAN as ablation in Table 3. We find that performance improves for almost all methods, further corroborating the effectiveness of our technique. In particular, only Llama-3 sees a marginal decline in performance.
Table 3: The ASR results and improvements at 1000 iterations. We note that functional homotopy is especially effective for gradient based methods such as GR and GCG. AutoDAN sees a marginal improvement for Llama-2, but a decline in Llama-3. We note that Llama-3 is very robust to AutoDAN in general.
ASR Improvement @ 1000 Iterations
Model GR $\rightarrow$ FH-GR GCG $\rightarrow$ FH-GCG AutoDAN $\rightarrow$ FH-AutoDAN
Llama-2 37.5 $\rightarrow$ 99.5 69.5 $\rightarrow$ 99.0 61.5 $\rightarrow$ 68.5
Llama-3 47.0 $\rightarrow$ 76.5 59.0 $\rightarrow$ 77.5 19.5 $\rightarrow$ 14.5
Appendix F Additional iteration distributions
@@
Appendix G FH level-set plot
Figure 6 presents a conceptual illustration of the homotopy optimization method, elucidating its underlying principles and operational dynamics. This visual representation provides a more detailed intuition of the method’s efficacy in navigating complex optimization landscapes.
Figure 6: Conceptual illustration of homotopy methods for suffix optimization. (a) Left: Greedy local search heuristic. The red region denotes successful suffixes. The search initiates from a starting point (black solid) and iteratively moves to the optimal neighboring input (dashed circle) based on loss values, potentially leading to local optima entrapment due to non-convexity. (b) Right: Homotopy approach. A series of progressively challenging optimization problems is constructed, with easier problems having larger solution spaces. The solution set gradually converges to that of the original problem. Adjacent problems in this continuum have proximal solutions, facilitating effective neighborhood search. Despite the underlying non-convexity, initiating from a near-optimal point simplifies each problem-solving step.
Appendix H Loss Convergence Analysis for GCG and FH-GR
In this section, we consider “hard” samples for Llama-2 and Llama-3 that GCG was unable to jailbreak, but FH-GR (initialized from checkpoint-500) was successful. Recall that both GCG and GR utilize identical loss objectives $F_{0}$, while FH-GR employs a sequence of loss objectives $(F_{n},...,F_{0})$, culminating in the same final loss as GCG. We designate $(F_{n},...,F_{0})$ as the homotopy loss, with input generation guided by $F_{i}$.
Figure 7 shows the change in average homotopy loss with iterations throughout the progression of homotopy. We find that easier optimization problems and the solution of the preceeding problem enables a consistently lower loss throughout the homotopy process.
Figure 8 examines how the adversarial strings found on weaker models by the functional homotopy, affect the average loss $F_{0}$ on the base model. We see that the loss consistently decreases, indicating that we are able to avoid local optima and successfully jailbreak the model faster.
Figure 9 shows the robustness of Functional Homotopy. Despite initializing the attack with stronger checkpoints (i.e., models fine-tuned for fewer epochs), we still find that loss converges more quickly that GCG and results in a jailbroken response from the model.
(a) FH-GR loss for Llama-2, across checkpoints
(b) FH-GR loss for Llama-3, across checkpoints
Figure 7: A loss comparison of GCG and FH-GR on “hard” samples. The red FH-GR Loss indicates the homotopy loss. We see that homotopy starts off with a substantially smaller loss, due to the misalignment process. We initialize FH-GR from checkpoint 500. As we iterate and successfully jailbreak an intermediate model, we replace it (as described in Algorithm 1), until we reach the base model by iteration 1000. We further note in Section 5 that FH-GR converges more quickly than GCG.
(a) FH-GR loss for base Llama-2
(b) FH-GR loss for base Llama-3
Figure 8: A loss comparison of GCG and FH-GR on “hard” samples. Unlike Figure 7, we look at the usefulness of the adversarial strings found by the functional homotopy, by applying them to the base model and computing the base GCG loss. The red FH-GR Loss indicates the base model loss objective value induced by the inputs found by the FH method. The loss decreases more consistently, before converging to a lower value overall and successfully jailbreaking the model compared to GCG.
(a) FH-GR loss for Llama-2, across checkpoints
(b) FH-GR loss for Llama-3, across checkpoints
[…]